OpenStudy (anonymous):

HELP ME! I cant get past this! 3(sqrt)2 - 2(sqrt)3 over 3(sqrt)2 + 2(sqrt)3

OpenStudy (anonymous):

\[3\sqrt{2}-2\sqrt{3}\div3\sqrt{2}+2\sqrt{3}\]

OpenStudy (anonymous):

\[\frac{ 3\sqrt2-2\sqrt3 }{ 3\sqrt2+2\sqrt3 }\] like this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you multiply top and bottom by the conjugate of the bottom... \[\frac{ 3\sqrt2-2\sqrt3 }{ 3\sqrt2+2\sqrt3 }=\frac{ (3\sqrt2-2\sqrt3) }{ (3\sqrt2+2\sqrt3) }\cdot\frac{ (3\sqrt2-2\sqrt3) }{ (3\sqrt2-2\sqrt3) }\] work that out and you'll get rid of the square roots in the denominator.

OpenStudy (anonymous):

Thats the part that im stuck on, i dont really understand how to work them out.

OpenStudy (anonymous):

did you learn how to FOIL or distribute?

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

Yes, but im still struggling with it. I dont understand what to do with the square roots and how to combine the different ones.

OpenStudy (anonymous):

okay... \[\frac{ (3\sqrt2-2\sqrt3)(3\sqrt2-2\sqrt3) }{ (3\sqrt2+2\sqrt3)(3\sqrt2-2\sqrt3) }=\]\[\frac{3\cdot3\cdot \sqrt2 \cdot \sqrt2-2(2\cdot3\cdot \sqrt2 \cdot \sqrt3)+2\cdot 2\cdot \sqrt3 \cdot \sqrt3}{3\cdot3\cdot \sqrt2 \cdot \sqrt2+(2\cdot3\cdot \sqrt2 \cdot \sqrt3)-(2\cdot3\cdot \sqrt2 \cdot \sqrt3)+2\cdot 2\cdot \sqrt3 \cdot \sqrt3}\]\[=\frac{ 9\cdot 2-2(6\cdot \sqrt6)+4\cdot 3 }{ 9 \cdot 2-4 \cdot 3 }=\frac{ 18-12\cdot \sqrt6+12 }{ 18-12 }=\frac{ 30-12\cdot \sqrt6 }{ 6 }\] and you should be able to simplify from there.

OpenStudy (anonymous):

Okay thank you so much! I'll do that now

OpenStudy (anonymous):

do you understand how to combine them?

OpenStudy (anonymous):

yes. would the answer be \[5-2\sqrt{6}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes, but do you understand how to multiply out the factors to get there, now?

OpenStudy (anonymous):

More than i did before but not completely

OpenStudy (anonymous):

rememeber, \(\sqrt a\cdot\sqrt b=\sqrt{a\cdot b}\) and \(a\sqrt b =a\cdot\sqrt b\)

OpenStudy (anonymous):

take it slow and write everything out and as you do more you'll get quicker and see shortcuts.

OpenStudy (anonymous):

Okay :) thank you so much. I wouldnt have just passed that online lesseon with out you